OpenStudy (anonymous):

I need help with one last question, using distance. Find the point on the graph of f(x)=x^2 that is closest to the point (3,1/2)

OpenStudy (paxpolaris):

use distance formula ..to write distance between (x,x^2) and (3, 0.5) then find minimum by taking derivative

OpenStudy (anonymous):

got the basic set up -> sqrt (x-2^2+(y-1/2)^2

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

one sec, so it would be sqrt (x-3)^2+(y-1/2)^2

OpenStudy (anonymous):

f'(x)=2x

OpenStudy (paxpolaris):

replace y with x^2 in distance formula

OpenStudy (anonymous):

okay, done

OpenStudy (anonymous):

so that would be sqrt(x-3)^2+(x^2-1/2)^2

OpenStudy (paxpolaris):

the set the derivative of the Distance formula to 0 ... to find minimum distance

OpenStudy (anonymous):

what is the y' of the distance formula. I'm confused

OpenStudy (paxpolaris):

note: to make it easier.. you can get rid of the sqrt ... because the sqrt is at minimum when what's inside the sqrt is at minimum

OpenStudy (anonymous):

good point work with the square of the distance, not the distance

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

but what would the derivative be?

OpenStudy (anonymous):

would it be 2(x-3)+2(x^2-1/2) ?

OpenStudy (anonymous):

\(D^2=(x-3)^2+(x^2-\frac{1}{2})^2\) multiply this mess out, combine like terms, take the derivative which should be easy enough because this is a polynomial

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

gotta do the algebra first

OpenStudy (paxpolaris):

gtg... my computer dying

OpenStudy (anonymous):

i mean not really, but you are going to need to find the zeros of your derivative, so you will have to multiply out in any case

OpenStudy (anonymous):

right now i hvae x^2+9+x^4+1/4

OpenStudy (anonymous):

have*

OpenStudy (anonymous):

okay, thank you so far for you help(:

OpenStudy (anonymous):

oh no that is not it

OpenStudy (anonymous):

i mean that is not what you get when you multiply out you get this http://www.wolframalpha.com/input/?i=%28x-3%29^2%2B%28x^2-1%2F2%29^2

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

How would I simplify that?

OpenStudy (anonymous):

oh , the y' is right there. okay. one sec

OpenStudy (anonymous):

4x^3-6

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

so would i set that to zero and solve?

OpenStudy (anonymous):

yeah wolfram gave away the store on this one you have the expression multiplied out and the derivative all in one

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

you even have the minimum there for you \[4x^3-6=0\] \[x^3=\frac{3}{2}\] \[x=\sqrt[3]{\frac{3}{2}}\]

OpenStudy (anonymous):

So that is the point closest to the point then. I'm looking at the process lol

OpenStudy (anonymous):

thanks a bunch (: